# HMT-Grasp: A Hybrid Mamba-Transformer Approach for Robot Grasping in Cluttered Environments

### v1

@@ IV-E Simulation and Real Robot Experiments
In this section, we first conducted robot grasping experiments in the Gazebo using models.First, we performed 10 isolated experiments for each object by randomly spawning it within the robot’s workspace. The results, presented in Table VI, demonstrate the effectiveness of our method in grasping various objects. The reasons for failures can be attributed to the complexity of object shapes and material properties, such as sponges being soft and deformable, making it difficult for the grasp detection model to predict an optimal grasp configuration. As illustrated in Fig. 3, we also conducted 10 rounds of object grasping in cluttered scenarios for both simulation and real robots. In the case of simulation experiments, the robot was able to outperform the selected state-of-the-art with a 7% margin (65% vs 72%). The model’s ability to balance local and global feature extraction, thanks to the Mamba blocks, contributed significantly to its improved performance in complex scenarios. The real robot achieved a success rate comparable to the simulation results, with 67% successful grasps. The failures in real-world experiments were primarily due to object occlusion. In particular, many objects are partially occluded by others, leading to attempts to grasp objects at an inappropriate angle or position.
@@
| Objects | mug | scissor | tea | banana | flash |
@@ V Conclusion
This paper presents a hybrid Mamba-Transformer model for robotic grasping in complex scenarios. By combining Mamba, Transformers, and CNNs, our method overcomes the limitations of ViTs and CNNs in visual grasp detection. We conducted extensive evaluations on the Cornell, Jacquard, and OCID-Grasp datasets, where our model consistently outperformed state-of-the-art methods, achieving an accuracy of 99.5% on the Cornell dataset. These results demonstrate the superior adaptability and precision of our approach. Additionally, real-world robotic experiments further validated the method’s effectiveness in dynamic and complex environments, highlighting its potential for practical applications. Future work could focus on incorporating real-time feedback mechanisms to dynamically adjust grasp plans during execution.

### v2

@@ IV-E Simulation and Real Robot Experiments
In this section, we first conducted robot grasping experiments in the Gazebo using models, as shown in Fig. 5. First, we performed 20 isolated experiments for each object by randomly spawning it within the robot’s workspace. The results, presented in Table VI, demonstrate the effectiveness of our method in grasping various objects. The results show that HMT-Grasp generally outperforms other methods across most objects. It achieved 100% success on the pear, surpassing GR-ConvNet (80%) and TF-Grasp (85%). It also performed well on can_coke (95%), exceeding all other methods. Although HMT-Grasp generally performed better, there were cases where the improvement was marginal. For example, HMT-Grasp and GR-ConvNet both reached 85% on the banana, and the performance difference for the peach was only . The reasons for failures can be attributed to the complexity of object shapes, material and collision mesh properties. For example, objects like the mug had lower success rates for both methods. The challenge with grasping this object lies in the difficulty of detecting stable grasp points, as small surface areas reduce friction and stability.
We conducted real-robot experiments in lightly, moderately, and heavily cluttered environments with complex backgrounds, as illustrated in Fig. 4. In the lightly cluttered setting, we tested twenty groups of five randomized objects per method to ensure diverse grasping scenarios. In the moderately cluttered setting, we conducted ten groups of ten randomized objects per method to evaluate grasping performance across various backgrounds. In the highly cluttered setting, we ran five experimental groups, each aiming to include twenty distinct objects to assess robustness under dense clutter. We compute the success rate as the ratio of successful grasps to the total number of grasp attempts. The results are presented in Table VII. Overall, our approach consistently showed better performance compared to the baseline methods. However, accuracy decreased from lightly to heavily cluttered scenarios for two primary reasons: (1) insufficient friction between the gripper jaws and certain smooth-surfaced objects, resulting in grasp failures; and (2) the inaccurate grasp prediction led to contacting the gripper with the neighboring objects, resulting in a grasp failure.
@@
| Objects | mug | scissor | tea_box | banana | flash |
@@
| Models | Success Rate () | ||
| Light Clutter | Moderate Clutter | High Clutter | |
| (5 Objects) | (10 Objects) | (20 Objects) | |
| GR-ConvNet [8] | 80 | 74 | 72 |
| TF-Grasp [9] | 82 | 75 | 70 |
| HMT-Grasp | 88 | 82 | 79 |
@@ V Conclusion
This paper presents a hybrid Mamba-Transformer model for robotic grasping in complex scenarios. By combining Mamba, Transformers, and CNNs, our method improves the performance of ViTs and CNNs in visual grasp detection. We conducted extensive evaluations on the Cornell, Jacquard, and OCID-Grasp datasets, where our model consistently outperformed state-of-the-art methods, achieving an accuracy of 99.5% on the Cornell dataset. These results demonstrate the superior adaptability and precision of our approach. Additionally, real-world robotic experiments further validated the method’s effectiveness in complex environments, highlighting its potential for practical applications. Future work could focus on incorporating real-time feedback mechanisms to adjust grasp plans dynamically during execution.